the resultant of two forces is 250n and the same are inclined at 30 n 45 with resultant,one on either side.calculate the magnitute of two forces.
 

Grade:12th pass

2 Answers

Manas Shukla
7 years ago
For the sake of simplicity lets assume
Manas Shukla
7 years ago
sorry forum bug...
Let the two forces be F1 and F2 and resultant be R.
Lets assume F1 lies on x axis So if F1 is on x axis , F2 will make 75degrees with x axis and resultant will make 30degrees with x axis.
So balancing the force components we get
R\cos 30 = F2 \cos 75 +F1   and
R\sin 30 = F2 \sin 75
Remember we can find the value of sin75 and cos75 by applying formula of sin(a+b) and cos(a+b). a =45 and b = 30
Solving we get F2 = 129.4N
substituting value of F2 in first equation
F1 = 183N